Calculate for the specific heat of 28.0 g of of mercury (Hg)

when 137 J of energy is involved when raised at an initial

temperature of 22.0°C to a final temperature of 57. 0C.


1
Expert's answer
2022-05-26T11:48:05-0400

q = m × C × ΔT

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C−1 g−1)

ΔT = change in temperature (°C)


q = 137 J

m = 28.0 g

C = unknown

ΔT = Tf − Ti = 57.0°C − 22.0°C = 35.0°C


Thus:

(137 J) = (28.0 g) × C × (35.0°C)

C = (137 J) / (28.0 g × 35.0°C) = 0.13979 J °C−1 g−1 = 0.140 J °C−1 g−1

C = 0.140 J °C−1 g−1


Answer: The specific heat of mercury (Hg) is 0.140 J °C−1 g−1